### v3

@@ 1 Introduction
Multi-armed bandits (MABs) have garnered significant attention as they provide well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems. In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards.
@@
This paper focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases. In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with mean equal to the inner product of $\theta^{\star}$ and the chosen action. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total expected rewards achieved by the policy and the maximum achievable expected reward.
@@
Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). Another popular strategy is Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
In the linear bandit setting, two types of regret are considered. The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al. (2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to polylogarithmic factors.
@@
On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of the noise and the algorithm. In this frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al. (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound. However, for a frequentist variant of LinTS referred to as TS-Freq, which inflates the posterior distribution, Agrawal and Goyal (2013) and Abeille et al. (2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret. This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$. Recent work by Hamidi and Bayati (2020b) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved. For the Greedy algorithm in linear bandit problems, no general theoretical guarantee exists. Moreover, it is known that both LinTS and Greedy might perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
@@
Despite the theoretical gaps in performance, LinTS has shown strong empirical performance (Russo et al., 2018), indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently. Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021). Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al. (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency. This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify, in a data-driven fashion, problematic instances where LinTS and Greedy could potentially fail and apply a “course-correction” to ensure competitive frequentist regret bounds? And can this be achieved without compromising their impressive empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
@@
  • We develop a real-time geometric analysis technique for the $d$-dimensional confidence ellipsoid surrounding $\theta^{\star}$. This method is crucial for maximizing the use of historical data, advancing beyond methods that capture only limited information from the confidence ellipsoid, such as a single numerical value. Consequently, this facilitates a more precise “course-correction”.
@@
  • We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, instance-dependent frequentist regret bound for them. This bound is efficiently computable using data observed from previous decision epochs.
@@
  • We introduce course-corrected variants of LinTS and Greedy that achieve minimax optimal frequentist regret. These adaptations maintain most of the desirable characteristics of the original algorithms.
The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid. This is in contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their focus on summary statistics.
@@ 1.1 Other Related literature
We note that Valko et al. (2014); Kocák et al. (2014, 2020); Kocák and Garivier (2020) consider the spectral of graphs and study Spectral Bandits. In their work, the rewards of arms are represented as smooth functions on a graph, allowing the exploitation of a low-rank structure for improved algorithmic performance and a regret guarantee that is independent of the original number of actions. In contrast, our approach leverages the spectral properties of the sample covariance of actions, which is distinct from graph spectral analysis. Additionally, our research addresses the broader context of stochastic linear bandit problems without presupposing any low-rank structure.
The concepts in this paper bear similarities to existing research that investigates various exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018), as well as recent studies by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a) that focus on using observed data for reducing exploration. Nonetheless, our methodology and the specific use of data markedly differ. For example, Bastani et al. (2021) emphasizes the minimum eigenvalue of the sample covariance matrix, essentially a single-parameter summary of the observed data. On the other hand, Hamidi and Bayati (2020a) uses information from the one-dimensional confidence intervals surrounding the reward values. A more closely related paper, Hamidi and Bayati (2020b), uses spectral information to enhance the performance of TS in LB problems. They look at a single summary statistics of the confidence set that they call thinness coefficient, to decide if the posterior of TS should be inflated. In contrast, with these studies, our approach capitalizes on the full geometric details of the $d$-dimensional confidence ellipsoid, utilizing richer geometric information.
1.2 Organization
The structure of this paper is as follows: Section 2 introduces the notations and problem formulation. In Section 3-4, we present the POFUL family of algorithms and analyze their frequentist regret. Sections 5-6 detail our data-driven approach and meta-algorithm for course-correction. Section 7 discusses simulations using synthetic and real-world datasets. The paper concludes in Section 8 with a summary of findings. All technical proofs, further discussions, and supplementary materials are provided in the appendices.
endcomment
@@ 2.1 Assumptions
Assumption 3 (Noise)
@@ 2.2 Regularized Least Square and Confidence Ellipsoid
Here, we review useful frequentist tools developed in Abbasi-Yadkori et al. (2011).
Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards. In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator. Fix a regularization parameter $\lambda$, the sample covariance matrix, and the Regularized Least Square (RLS) estimate are respectively given by
$V_{t}=\lambda I+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
The following proposition in Abbasi-Yadkori et al. (2011) states that the RLS estimate $\widehat{\theta}_{t}$ concentrates around $\theta^{\star}$ properly.
@@
Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that $\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)$ and $\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$ for all $x\in\mathbb{R}^{d}$, where
$\beta_{t}^{RLS}(\delta)=R\sqrt{2\log\frac{(\lambda+t)^{d/2}\lambda^{-d/2}}{\delta}}+\sqrt{\lambda}S.$ (2)
Hereafter, we might omit the dependence of $\beta_{t}$ on $\delta$ if there is no ambiguity. Propostion 1 enables us to construct the following sequence of confidence ellipsoids.
@@
Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as $\mathcal{E}_{t}^{RLS}(\delta)=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)\}$.
The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret. Essentially, this proposition suggests that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
@@
If $\lambda>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda I)}\leq 2d\log\left(1+\frac{t}{\lambda}\right).$
@@ 3 POFUL Algorithms
In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). It allows us to analyze the frequentist regret of common algorithms in a unified way.
At a high level, POFUL is designed to encompass the exploration mechanism of OFUL and LinTS. POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
@@
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$,
$\mathbb{P}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]\geq 1-\delta^{\prime},$
where we define the pivot ellipsoid as $\mathcal{E}_{t}^{PVT}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t}(\delta)\right\}$.
@@
$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$ (3)
A pseudocode representation of POFUL is provided in Algorithm 1.
Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid. We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$. By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
@@
0:  $T$, $\delta$, $\lambda$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
@@
     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$
@@
OFUL
For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$ Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
LinTS and TS-Freq
Linear Thompson Sampling (LinTS) algorithm is a generic randomized algorithm that samples from a distribution constructed from the RLS estimate at each step. At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, and $\iota^{TS}_{t}$ is an inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability.
@@
Fix any $\delta\in(0,1)$. $\mathcal{D}^{SA}(\delta)$ is a distribution such that $\mathbb{P}_{\theta\sim\mathcal{D}^{SA}(\delta)}\left[\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq 1\right]\geq 1-\delta.$
LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$. We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm. On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), namely TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
Greedy
Greedy is a special case of POFUL where $\iota_{t}=\tau_{t}=0$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$ for all $t\in[T]$.

### v5

@@
Geometry-Aware Bandit Algorithms
[…]
Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits
[…]
This paper is motivated by recent research in the $d$-dimensional stochastic linear bandit literature, which has revealed an unsettling discrepancy: algorithms like Thompson sampling and Greedy demonstrate promising empirical performance, yet this contrasts with their pessimistic theoretical regret bounds. The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances. To address this, we propose a new data-driven technique that tracks the geometric properties of the uncertainty ellipsoid around the main problem parameter. This methodology enables us to formulate a data-driven frequentist regret bound, which incorporates the geometric information, for a broad class of base algorithms, including Greedy, OFUL, and Thompson sampling. This result allows us to identify and “course-correct” problem instances in which the base algorithms perform poorly. The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$ for a $T$-period decision-making scenario, effectively maintaining the desirable attributes of the base algorithms, including their empirical efficacy. We present simulation results to validate our findings using synthetic and real data.
[…]
Linear bandit, Thompson sampling, Greedy, Data-driven exploration
@@ 1 Introduction
Multi-armed bandits (MABs) provide a framework for studying the exploration-exploitation trade-off in sequential decision-making, where a decision-maker selects actions and observes uncertain rewards. This extends to contextual bandits with features or covariates, as shown in numerous applications (Langford and Zhang, 2008; Li et al., 2010; Tewari and Murphy, 2017; Zhou et al., 2020; Villar et al., 2015; Bastani and Bayati, 2020; Cohen et al., 2020). This paper focuses on a well-studied class of models that captures both MABs and contextual bandits as special cases while being amenable to theoretical analysis: the stochastic linear bandit (LB) problem. In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with a mean equal to the inner product of $\theta^{\star}$ and the chosen action. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total expected rewards achieved by the policy and the maximum achievable expected reward.
@@
Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the optimism principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). Another popular strategy is Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
@@
In the linear bandit setting, two regret types are considered. Bayesian regret treats parameter $\theta^{\star}$ as a random variable with a prior distribution, averaging regret over noise, algorithm randomness, and parameter randomness, measuring expected performance across parameter realizations. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the Thompson Sampling (TS) heuristic, referred to as LinTS, matching the minimax optimal bound by Dani et al. (2008). Here, $\widetilde{\mathcal{O}}$ denotes asymptotic order up to polylogarithmic factors. Frequentist regret assumes fixed unknown parameter $\theta^{\star}$, averaging only over noise and algorithm randomness. The OFUL algorithm (Abbasi-Yadkori et al., 2011) achieves an optimal $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound. However, TS-Freq, a frequentist LinTS variant with inflated posterior variance, only achieves $\widetilde{\mathcal{O}}(d\sqrt{dT})$ (Agrawal and Goyal, 2013; Abeille et al., 2017), suboptimal by factor $\sqrt{d}$. Hamidi and Bayati (2020b) confirms this inflation is necessary and LinTS’s frequentist regret cannot be improved. The Greedy algorithm lacks theoretical guarantees for linear bandit problems (Lattimore and Szepesvari, 2017), suggesting both LinTS and Greedy may perform suboptimally.
@@
Despite the theoretical gaps, LinTS demonstrates strong empirical performance (Russo et al., 2018), suggesting posterior distribution inflation may be unnecessary in most scenarios. Similarly, the Greedy algorithm performs well in typical cases (Bietti et al., 2021). While optimism-based algorithms are computationally expensive (generally NP-hard (Dani et al., 2008; Russo and Van Roy, 2014; Agrawal, 2019)), LinTS and Greedy maintain computational efficiency. This disparity between theoretical, computational, and empirical performance prompts two questions: Can we identify problematic instances for LinTS and Greedy in a data-driven way and apply ”course-correction” to ensure competitive frequentist regret bounds? Can this be achieved while preserving their empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
@@
1. We develop a real-time geometric analysis technique for the $d$-dimensional confidence ellipsoid surrounding $\theta^{\star}$. This method is crucial for maximizing the use of historical data, advancing beyond methods that capture only limited information from the confidence ellipsoid, such as a single numerical value. Consequently, this facilitates a more precise “course-correction”.
@@
2. We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, data-driven frequentist regret bound for them. This bound is efficiently computable using data observed from previous decision epochs.
@@
3. We introduce course-corrected variants of LinTS and Greedy that achieve minimax optimal frequentist regret. These adaptations maintain most of the desirable characteristics of the original algorithms.
@@ 1.1 Other Related Literature
Our work is closely related to three main research streams: methodological foundations of linear bandits, bandit algorithms utilizing spectral properties, and data-driven exploration techniques. While these works share some similarities with our approach, we highlight the key differences and the unique aspects of our methodology.
From a methodological perspective, our regret analysis builds upon the foundations laid by Abbasi-Yadkori et al. (2011), Agrawal and Goyal (2013), and Abeille et al. (2017). However, a key distinguishing factor is that our approach does not rely on optimistic samples, which is a departure from previous methods. This means that the algorithms we study do not always choose actions that are expected to perform better than the true optimal action. By allowing non-optimistic samples, we avoid the need to inflate the posterior distribution, a requirement in the works of Agrawal and Goyal (2013) and Abeille et al. (2017).
Our use of spectral information in bandit algorithms bears some resemblance to the study of Spectral Bandits (Valko et al., 2014; Kocák et al., 2014, 2020; Kocák and Garivier, 2020). These works represent arm rewards as smooth functions on a graph, leveraging low-rank structures to improve algorithmic performance and obtain regret guarantees independent of the number of actions. In contrast, our approach exploits the spectral properties of the action covariance matrix, which is distinct from graph spectral analysis. Moreover, our research tackles the broader context of stochastic linear bandits without assuming any low-rank structure.
Our work also shares conceptual similarities with research on exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018) and data-driven exploration reduction (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a). However, our methodology and data utilization differ significantly. For instance, Bastani et al. (2021) focuses on the minimum eigenvalue of the covariance matrix, a single-parameter summary of the observed data, while Hamidi and Bayati (2020a) uses information from one-dimensional reward confidence intervals. The work of Hamidi and Bayati (2020b) is more closely related to ours, as it employs spectral information to improve the performance of Thompson Sampling in linear bandits. They use a single summary statistic called the thinness coefficient to decide whether to inflate the posterior. In contrast, our approach leverages the full geometric details of the $d$-dimensional confidence ellipsoid, harnessing richer geometric information.
@@ 2 Setup and Preliminaries
Assumption 3 (Subgaussian reward noise)
@@ 2.1 Regularized Least Square and Confidence Ellipsoid
In this subsection, we review the useful frequentist tools developed by Abbasi-Yadkori et al. (2011) for estimating the unknown parameter $\theta^{*}$ in linear bandit (LB) problems.
Consider an arbitrary sequence of actions $(x_{1},\dots,x_{t})$ and their corresponding rewards $(r_{1},\dots,r_{t})$. In LB problems, the parameter $\theta^{*}$ is typically estimated using the regularized least squares (RLS) estimator. Let $\lambda_{\text{reg}}$ be a fixed regularization parameter. The sample covariance matrix $V_{t}$ and the RLS estimate $\widehat{\theta}_{t}$ are defined as follows:
$V_{t}=\lambda_{\text{reg}}I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
The following proposition from Abbasi-Yadkori et al. (2011) establishes that the RLS estimate $\widehat{\theta}_{t}$ concentrates around the true parameter $\theta^{*}$ with high probability.
@@
Let $\delta\in(0,1)$ be a fixed confidence level. Then, with probability at least $1-\delta$, it holds for all $x\in\mathbb{R}^{d}$ that
$\|\widehat{\theta}_{t}-\theta^{\star}\|_{V_{t}}\leq\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS},\quad|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle|\leq\|x\|_{V_{t}^{-1}}\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}$
where the confidence bound $\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}$ is defined as
$\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}=R\sqrt{2\log(\lambda_{\text{reg}}+t)^{d/2}\lambda_{\text{reg}}^{-d/2}{\delta}^{-1}}+\sqrt{\lambda_{\text{reg}}}S.$ (2)
Proposition 1 enables us to construct the following sequence of confidence ellipsoids.
@@
Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as
$\mathcal{E}_{t,\delta,\lambda_{\text{reg}}}^{RLS}=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t,\delta,\lambda_{\text{reg}}}^{RLS}\}\,.$
The next proposition, known as the elliptical potential lemma, plays a central role in bounding the regret. This proposition provides the key element in the work of Abbasi-Yadkori et al. (2011), showing that the cumulative prediction error incurred by the action sequence used to estimate $\theta^{*}$ is small.
@@
If $\lambda_{\text{reg}}>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda_{\text{reg}}I)}\leq 2d\log(1+\frac{t}{\lambda_{\text{reg}}})\,.$
@@ 3 POFUL Algorithms
In this section, we introduce POFUL (Pivot OFUL), a generalized framework of OFUL. This framework enables a unified analysis of frequentist regret for common algorithms.
At a high level, POFUL is designed to encompass the exploration mechanism of OFUL and LinTS. POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This is formalized in the next definition.
@@
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$, $\mathbb{P}[\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}]\geq 1-\delta^{\prime}$, where we define the “pivot ellipsoid” as $\mathcal{E}_{t,\delta,\lambda_{\text{reg}}}^{PVT}\coloneqq\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t,\delta,\lambda_{\text{reg}}}\}$.
@@
$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\,,$ (3)
as shown in a pseudocode representation in Algorithm 1 and illustrated in Figure 1(a).
Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid. We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$.
@@
0:  $T$, $\delta$, $\lambda_{\text{reg}}$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
  Initialize $V_{0}\leftarrow\lambda_{\text{reg}}I_{d}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
@@
     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$
@@
[…]
Figure 1: (a) POFUL algorithms illustration for general $\iota_{t}$ and $\tau_{t}$. (b) Special cases: OFUL ($\iota_{t}=0$, $\tau_{t}=1$), TS ($\tau_{t}=0$), and Greedy ($\iota_{t}=\tau_{t}=0$).
We demonstrate that POFUL encompasses OFUL, LinTS, TS-Freq, and Greedy as special cases, as illustrated in Figure 1(b).
Example 1 (OFUL)
For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$. Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
Before describing how TS can be derived as an instance of POFUL, we introduce a definition.
@@
Let $\delta\in(0,1)$. We define $\mathcal{D}^{SA}(\delta)$ as a distribution satisfying $\mathbb{P}_{\eta\sim\mathcal{D}^{SA}(\delta)}\left[\|\eta\|\leq 1\right]\geq 1-\delta$.
Example 2 (TS)
Linear Thompson Sampling (LinTS) algorithm is a generic randomized algorithm that samples from a distribution constructed from the RLS estimate at each step. At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-1/2}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, $\iota^{TS}_{t}$ is inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability. LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-1/2}\eta_{t}$. Setting $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm, while setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), namely TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
Example 3 (Greedy)
Greedy is a special case of POFUL with $\iota_{t}=\tau_{t}=0$, $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, $\forall t$.